OpenStudy (winsteria):

MEDAL AND FAN What is the solution to the equation? 2(x ��������� 7) = 10x + 18 A. x = -15/8 B. x= -4 C. x= 15/8 D. x= 4

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

x = –4 x = - 15/8 x = 15/8 x = 4

OpenStudy (winsteria):

Thank you

OpenStudy (anonymous):

np.if you need more help just tag me and i will help you:)